OpenStudy (anonymous):

tan(x/2)+2sin2x=cscx

OpenStudy (jdoe0001):

\(\bf tan\left(\frac{x}{2}\right)+2sin(2x)=csc(x)\quad ?\qquad tan\left(\frac{x}{2}\right)+2sin^2(x)=csc(x)\quad ?\)

OpenStudy (anonymous):

The first one

OpenStudy (jdoe0001):

what are you meant to do with it?